# Bézier Splatting for Fast and Differentiable Vector Graphics Rendering

### v3

@@ 4.1 Experimental Setups
Implementation details. We implement Bézier Splatting in PyTorch [19] and optimize it by using the Adam optimizer [12] with a StepLR learning rate scheduler. The learning rate is initialized at 0.01 for color, 2e-4 for Bézier curve control points, and 0.1 for opacity. For the pruning and densification strategy, the opacity threshold is set to 0.02, and the overlap threshold based on Axis-Aligned Bounding Boxes (AABB) is set to 0.9. Following the approach in LIVE [17], new curves are initialized in a circular pattern, and the number of added curves matches the number of removed ones to maintain a constant total curve count. Open curves are optimized for 15,000 iterations, and closed curves are optimized for 10,000 iterations. Pruning and densification are applied every 400 steps until the last but 1,000 steps, after which they are halted to stabilize the representation.
Datasets. We comprehensively evaluate our method across different image domains. We use the publicly available DIV2K [23] dataset for evaluating natural images. Due to the high computational cost of baseline method LIVE [17], we uniformly subsample the DIV2K dataset by selecting one out of every four images, resulting in a final evaluation set of 200 images from the original 800-image dataset. Additionally, we test our method on non-photorealistic images, including the artwork images from Clipart1K dataset [9] and cartoon images from Danbooregions dataset [31], to demonstrate its effectiveness for diverse types of images, as shown in Fig. 4.
Table 2: Quantitative evaluation and optimization efficiency of differentiable VG methods on the DIV2K dataset [23].
@@
[…]
Open DiffVG [14] 0.552 19.83 0.563 18.9min 0.587 21.47 0.537 22.0min 0.616 22.62 0.517 30.6min
@@
Closed DiffVG [14] 0.578 20.69 0.548 16.4min 0.601 21.82 0.531 18.5min 0.631 22.95 0.509 25.1min
LIVE [17] 0.576 20.09 0.543 2.6h 0.611 21.70 0.521 4.2h 0.648 23.11 0.495 5.1h
LIVSS [26] 0.586 17.71 0.542 39.2min 0.630 18.71 0.530 54.3min 0.678 19.83 0.517 1.4h
@@ 4.2 Method Comparison
Table 1 reports the forward and backward runtime for processing 2,048 curves on an image with a resolution of 2,040$\times$1,344. Compared to DiffVG [14], our method significantly accelerates VG rasterization by 31.4$\times$ faster per forward step and 149.2$\times$ faster per backward step for open curves. For closed curves, our color filling strategy requires sampling 20 additional Bézier curves per closed curve, while our method remains highly efficient, achieving 6$\times$ faster forward and 18.2$\times$ faster backward computation.
Table 2 quantitatively evaluates the quality of differentiable VG representations by three commonly used metrics, MS-SSIM [25], PSNR, and LPIPS [33]. Our method demonstrates higher optimization efficiency and rendering fidelity for both open and closed curves with different curve numbers.
Fig. 3 visually compares DiffVG [14], LIVE [3], [26], and our Bézier Splatting under 512 closed curves and 2048 open curves. Compared to DiffVG [14], our method captures significantly more fine-grained textures. Compared to LIVE [17], our method effectively enhances rendering quality, resulting in higher-fidelity results. This improvement stems from our method’s ability to globally optimize all curves, rather than LIVE’s layer-by-layer path-adding strategy. Compared to LIVSS [26], our method demonstrates superior ability in preserving fine structural details, especially in text regions. Due to its heavy reliance on layer-wise semantic simplification, LIVSS struggles to optimize regions where semantic information is ambiguous or difficult to extract. Moreover, by optimizing the entire VG simultaneously, our approach ensures a cohesive and natural appearance, avoiding the accumulation of errors and inconsistencies introduced by layer-wise updates.
As shown in Fig. 4, Bézier Splatting consistently renders high-quality VGs across various image domains from photorealistic natural images, watercolor paintings, to cartoon images. Unlike learning-based image vectorization methods [15, 20, 1], which can not easily generalize to out-of-domain data, our method can flexibly handle different types of images.
Figure 5: As a differentiable VGs renderer, Bézier Splatting can integrate with topology-aware strategies such as layer-wise vectorization [17] to further improve compositionality and detail.
@@
4.3 Adaptability with Layer-wised Image Vectorization
Existing works such as LIVE [17], SGLIVE [36] and LIVSS [26] have demonstrated that layer-wise vectorization strategies can significantly improve the topology and compositionality of DiffVG-generated vector graphics [14]. Since the proposed Bézier Splatting is a fast and differentiable VGs renderer alternative to DiffVG, is naturally compatible with these topology-aware strategies and can leverage them to further enhance the structural quality of the optimized vector outputs. As shown in Fig. 5, we incorporate the layer-wise vectorization approach from LIVE [17] into the Bézier Splatting optimization process This integration yields improved compositional integrity and finer details for regions with complex patterns such as the butterfly’s wings.
@@ 5 Conclusion
This work has presented Bézier Splatting, a novel differentiable vector graphics (VGs) representation that leverages Gaussian splatting for efficient Bézier curve optimization. Our method achieves 30$\times$ faster forward computation and 150$\times$ faster backward computation in rasterization compared to existing methods, while also delivering high rendering fidelity. Additionally, our adaptive pruning and densification strategy improves optimization by dynamically adjusting curve placement during optimization. Extensive experiments have demonstrated that Bézier Splatting outperforms existing differentiable VG methods in both training efficiency and visual quality, making it a promising solution for scalable applications of VGs.
Limitation and future work. The closed curves in Bézier Splatting are enforced to have convex shapes using Xing loss [17] to prevent false interpolation results, which may slightly reduce the capacity of VG representations. Additionally, the closed curves require sampling more Gaussian points to represent area boundaries precisely, leading to slower computation compared to open curves. An interesting future direction is to leverage the efficiency and differentiability of our approach for high-quality VG synthesis applications such as text-to-VG generation or text-based VG editing.
@@ Appendix A Ablation Study
The effectiveness of adaptive pruning and densification and layer-wise vectorization strategy. We conduct an ablation study to evaluate the effectiveness of our adaptive pruning and densification strategy. Quantitative results are shown in Table 3. To further investigate the flexibility of our framework, we incorporate a layer-wise curve addition strategy inspired by LIVE [17], where curves are progressively added during training. Although this strategy achieves slightly better reconstruction metrics, it doubles the training time compared to our adaptive pruning and densification approach, making it less favorable for time-sensitive applications. Fig. 7 shows our result with layer-wise image vectorization strategy.
Table 3: An ablation study on adaptive pruning and densification and layer-wise training strategy, evaluated on 512 closed curves with DIV2K dataset [23].
@@
Figure 7: Our Bézier Splatting is fully compatible with the layer-wise vectorization strategies [17].
@@
Appendix B Convert Bézier Splatting to standard SVG Format
As described in Algorithm 1, we convert our optimized Bézier Splatting representation into a standard SVG file for compatibility with downstream vector editing tools. Each curve is represented by a set of $13k+$ 1 control points $\mathcal{C}_{i}\in\mathbb{R}^{(3k+1)\times 2}$, corresponding to $k$ continuous cubic Bézier segments. These control points are normalized to the range $[-1,1]$ and are first transformed into pixel coordinates according to the canvas size $(W,H)$.
@@ Appendix B Convert Bézier Splatting to standard SVG Format
Algorithm 1 Convert Bézier Splatting to Standard SVG.
@@
Appendix C Bézier Splatting Supports Flexible Attributes of Curves
All 2D Gaussians in our method are generated via a differentiable sampling algorithm, which makes it straightforward to incorporate user-defined shape attributes, such as linear-gradient fills in color or opacity. To demonstrate the extensibility of our model, we evaluate the multi-opacity scheme for open curves: each Bézier segment within an open curve is assigned an independent opacity value. This design enables users to either maintain consistent sampling point appearance across segments or apply customized interpolation strategies. This strategy significantly improves vectorization quality for open curves, as shown in Table  5. Since closed curves are the primary representation format in SVG, we adopt a uniform setting for opacity and color in those regions to ensure a full compatibility with existing vector graphic tools. Nonetheless, users can still follow the above multi-opacity scheme to implement linear-gradient fills for both opacity and color, enhancing the flexibility and expressiveness of vector graphics.
Table 5: An ablation study on the number of opacity parameters per curve, evaluated on 512 open curves with DIV2K dataset [23].
@@ Appendix C Bézier Splatting Supports Flexible Attributes of Curves
Appendix D More Comparisons
We present more qualitative comparisons on the DIV2K [23] dataset. As shown in Fig. 9, Fig. 10, Fig. 11, Fig. 12, and Fig. 13, our method shows consistent improvements over the baselines, including better preservation of fine details, higher rendering fidelity, fewer visual artifacts, and more accurate geometric structures. Furthermore, as the number of curves increases, our method demonstrates a significantly improved representational ability, enabling more precise reconstruction of complex shapes, and fine-grained structures.
We further evaluate our method on another natural image dataset, Kodak [13]. Due to the slow vectorization speed of LIVE [17], we compare with DiffVG on open curves only. As shown in Table 6, our method consistently outperforms DiffVG across quantitative metrics including PSNR, SSIM, and LPIPS.
Table 6: A quantitative evaluation on the Kodak dataset [13] with 256 to 1024 curves. We report results on open curves.
@@ Appendix D More Comparisons
Appendix E More Image Vectorization Results
Fig. 14, Fig. 15, and Fig. 16 show more results on natural images from DIV2K [23] and Kodak [13], as well as animation images from DanbooRegion [31], respectively. The results demonstrate that both the global structure and local texture of the images are well reconstructed, highlighting the effectiveness of our approach in capturing fine details and complex shapes.
Figure 9: A qualitative comparison of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 10: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 11: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 12: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 13: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 14: More image vectorization results by Bézier Splatting on DIV2K dataset [23].
Figure 15: More image vectorization results by Bézier Splatting on Kodak [13] dataset.
Figure 16: More image vectorization results by Bézier Splatting on DanbooRegion dataset [31].

### v4

@@ 4.1 Experimental Setups
Implementation details. We implement Bézier Splatting in PyTorch [paszke2019pytorch] and optimize it by using the Adam optimizer [2015-kingma] with a StepLR learning rate scheduler. The learning rate is initialized at 0.01 for color, 2e-4 for Bézier curve control points, and 0.1 for opacity. For the pruning and densification strategy, the opacity threshold is set to 0.02, and the overlap threshold based on Axis-Aligned Bounding Boxes (AABB) is set to 0.9. Following the approach in LIVE [xu2022live], new curves are initialized in a circular pattern, and the number of added curves matches the number of removed ones to maintain a constant total curve count. Open curves are optimized for 15,000 iterations, and closed curves are optimized for 10,000 iterations. Pruning and densification are applied every 400 steps until the last but 1,000 steps, after which they are halted to stabilize the representation.
Datasets. We comprehensively evaluate our method across different image domains. We use the publicly available DIV2K [Timofte_2017_CVPR_Workshops] dataset for evaluating natural images. Due to the high computational cost of baseline method LIVE [xu2022live], we uniformly subsample the DIV2K dataset by selecting one out of every four images, resulting in a final evaluation set of 200 images from the original 800-image dataset. Additionally, we test our method on non-photorealistic images, including the artwork images from Clipart1K dataset [inoue2018cross] and cartoon images from Danbooregions dataset [DanbooRegion2020], to demonstrate its effectiveness for diverse types of images, as shown in Fig. 4.
Table 2: Quantitative evaluation and optimization efficiency of differentiable VG methods on the DIV2K dataset [Timofte_2017_CVPR_Workshops].
@@
$SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $\text{Opt.}^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $\text{Opt.}^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $\text{Opt.}^{\downarrow}$
Open DiffVG [Li:2020:DVG] 0.552 19.83 0.563 18.9min 0.587 21.47 0.537 22.0min 0.616 22.62 0.517 30.6min
@@
Closed DiffVG [Li:2020:DVG] 0.578 20.69 0.548 16.4min 0.601 21.82 0.531 18.5min 0.631 22.95 0.509 25.1min
LIVE [xu2022live] 0.576 20.09 0.543 2.6h 0.611 21.70 0.521 4.2h 0.648 23.11 0.495 5.1h
LIVSS [livss] 0.586 17.71 0.542 39.2min 0.630 18.71 0.530 54.3min 0.678 19.83 0.517 1.4h
@@ 4.2 Method Comparison
[…]
Table 2 quantitatively evaluates the quality of differentiable VG representations by three commonly used metrics, MS-SSIM [583d8b59c51a4f3e9bd0b90731111848], PSNR, and LPIPS [zhang2018perceptual]. Our method demonstrates higher optimization efficiency and rendering fidelity for both open and closed curves with different curve numbers.
Fig. 3 visually compares DiffVG [Li:2020:DVG], LIVE [Du:2023:IVE], [livss], and our Bézier Splatting under 512 closed curves and 2048 open curves. Compared to DiffVG [Li:2020:DVG], our method captures significantly more fine-grained textures. Compared to LIVE [xu2022live], our method effectively enhances rendering quality, resulting in higher-fidelity results. This improvement stems from our method’s ability to globally optimize all curves, rather than LIVE’s layer-by-layer path-adding strategy. Compared to LIVSS [livss], our method demonstrates superior ability in preserving fine structural details, especially in text regions. Due to its heavy reliance on layer-wise semantic simplification, LIVSS struggles to optimize regions where semantic information is ambiguous or difficult to extract. Moreover, by optimizing the entire VG simultaneously, our approach ensures a cohesive and natural appearance, avoiding the accumulation of errors and inconsistencies introduced by layer-wise updates.
As shown in Fig. 4, Bézier Splatting consistently renders high-quality VGs across various image domains from photorealistic natural images, watercolor paintings, to cartoon images. Unlike learning-based image vectorization methods [lopes2019learned, reddy2021im2vec, Cao_2023_CVPR], which can not easily generalize to out-of-domain data, our method can flexibly handle different types of images.
Figure 5: As a differentiable VGs renderer, Bézier Splatting can integrate with topology-aware strategies such as layer-wise vectorization [xu2022live] to further improve compositionality and detail.
@@
4.3 Layer-wise Image Vectorization
Existing works such as LIVE [xu2022live], SGLIVE [zhou2024segmentation] and LIVSS [livss] have demonstrated that layer-wise vectorization strategies can significantly improve the topology and compositionality of DiffVG-generated vector graphics [Li:2020:DVG]. Since the proposed Bézier Splatting is a fast and differentiable VGs renderer alternative to DiffVG, is naturally compatible with these topology-aware strategies and can leverage them to further enhance the structural quality of the optimized vector outputs. As shown in Fig. 5, we incorporate the layer-wise vectorization approach from LIVE [xu2022live] into the Bézier Splatting optimization process This integration yields improved compositional integrity and finer details for regions with complex patterns such as the butterfly’s wings.
@@ 5 Conclusion
[…]
Limitation and future work. The closed curves in Bézier Splatting are enforced to have convex shapes using Xing loss [xu2022live] to prevent false interpolation results, which may slightly reduce the capacity of VG representations. Additionally, the closed curves require sampling more Gaussian points to represent area boundaries precisely, leading to slower computation compared to open curves. An interesting future direction is to leverage the efficiency and differentiability of our approach for high-quality VG synthesis applications such as text-to-VG generation or text-guided VG editing.
6 Acknowledgement
This work was supported in part by the National Science Foundation (NSF) and SC EPSCoR Program under award number #OIA-2242812. This research used in part resources on the Palmetto Cluster at Clemson University under NSF awards MRI 1228312, II NEW 1405767, MRI 1725573, and MRI 2018069. The views expressed in this article do not necessarily represent the views of NSF or the United States government. The authors also thank Dr. Yiqi Zhong and Dr. Luming Liang for beneficial discussions.
@@ Appendix A Ablation Study
The effectiveness of adaptive pruning and densification and layer-wise vectorization strategy. We conduct an ablation study to evaluate the effectiveness of our adaptive pruning and densification strategy. Quantitative results are shown in Table 3. To further investigate the flexibility of our framework, we incorporate a layer-wise curve addition strategy inspired by LIVE [xu2022live], where curves are progressively added during training. Although this strategy achieves slightly better reconstruction metrics, it doubles the training time compared to our adaptive pruning and densification approach, making it less favorable for time-sensitive applications. Fig. 7 shows our result with layer-wise image vectorization strategy.
Table 3: An ablation study on adaptive pruning and densification and layer-wise training strategy, evaluated on 512 closed curves with DIV2K dataset [Timofte_2017_CVPR_Workshops].
@@
Figure 7: Our Bézier Splatting is fully compatible with the layer-wise vectorization strategies [xu2022live].
@@
Table 5: A quantitative comparison between Adobe Image Trace [illustrator2025] and our method across different images.
Image ID 0004 0008 0012 0016 0020
Method Adobe [illustrator2025] Ours Adobe [illustrator2025] Ours Adobe [illustrator2025] Ours Adobe [illustrator2025] Ours Adobe [illustrator2025] Ours
$SSIM^{\uparrow}$ 0.835 0.849 0.645 0.637 0.658 0.658 0.627 0.628 0.814 0.822
$PSNR^{\uparrow}$ 26.43 28.62 23.26 24.33 22.29 22.87 23.87 24.54 26.67 28.31
$LPIPS^{\uparrow}$ 0.380 0.387 0.489 0.489 0.523 0.518 0.534 0.543 0.505 0.503
Comparison with conventional image vectorization methods. We conduct a comparison against the Image Trace from Adobe Illustrator [illustrator2025] on five images from the DIV2K dataset. Both methods are evaluated using the same number of parameters (around 20K). As shown in Table 5, our method achieves higher fidelity, with an average PSNR of 25.734 db compared to 24.904 db of Image Trace.
Arbitrary resolution rendering. Our Bézier Splatting representation naturally supports rendering at arbitrary resolutions because all Gaussian parameters are analytically derived from the underlying Bézier curves. When higher-resolution images are required, proportionally increasing the sampling rate is sufficient to maintain reconstruction fidelity. For example, doubling the resolution simply doubles the sampling density. As shown in Table 6, rendering at higher resolutions (4K or 8K) with increased sampling preserves a comparable level of visual quality to the 2K baseline, without introducing noticeable artifacts. We report results on the first four images from the [Timofte_2017_CVPR_Workshops] to illustrate this property.
[…]
Image ID 2K (orig.) 4K 8K
orig. ($\times$2S) orig. ($\times$4S)
0004 26.8976 27.1553 27.2761 25.9800 26.6594
0008 21.2063 21.0614 21.4340 19.8318 21.1214
0012 19.8816 19.9079 20.0900 18.9389 19.7760
0016 22.3456 22.4327 22.5408 21.7088 22.3140
Appendix B Details of Adaptive Pruning and Densification
To provide a clearer understanding of our optimization behavior, we present here the detailed procedure of the pruning–densification algorithm 1 used in our Bézier-splatting framework.This algorithm is designed to mitigate the local-minima issue observed in 3D Gaussian Splatting (3DGS) and is adapted in our Bézier-splatting framework, where suboptimal primitive initialization leaves certain regions insufficiently covered. Inspired by the pruning–splitting mechanism of 3DGS [1], our method dynamically reallocates Bézier curves according to the reconstruction error map: curves with low opacity, small area, or high overlap with nearby curves of similar color are pruned, while new curves are inserted into regions with high reconstruction error to improve coverage. This reallocation keeps the total curve budget fixed, enhances reconstruction quality, and helps the optimization escape poor local minima. For open curves, we additionally introduce a splitting rule: if the middle opacity of a curve is more than 0.5 lower than that of both endpoints, the curve is split into two segments to better adapt to local structure variations.
Algorithm 1 Adaptive Pruning and Densification for Closed Bézier Curves
Notation:
  $\mathbf{opacity}(b)$ — mean opacity of curve $b$.
  $\mathbf{area}(b)$ — area enclosed by curve $b$.
  $\mathbf{colordiff}(b_{i},b_{j})$ — Euclidean color difference between curves $b_{i}$ and $b_{j}$.
  $\mathbf{IoU}(b_{i})$ — intersection-over-union between the bounding box of $b_{i}$ and neighboring curves with $\mathbf{colordiff}<0.03$.
  $\mathbf{ConnectedComponents}(E)$ — connected error regions extracted from the quantized error map.
Input : Closed Bézier curve set $\mathcal{B}=\{b_{1},\dots,b_{N}\}$; opacity map $\alpha$; error map $E$; iteration index $t$.
Output: Updated curve set $\mathcal{B}^{\prime}$.
 Initialize $\mathcal{B}^{\prime}\leftarrow\mathcal{B}$.
Pruning phase
for $i\leftarrow 1$ to NN do
  if $\mathbf{opacity}(b_{i})<\tau_{\mathbf{opacity}}(t)$ then
   Remove $b_{i}$ from $\mathcal{B}^{\prime}$. {Low-opacity pruning}
  end if
  if $\mathbf{area}(b_{i})<\tau_{\mathbf{area}}$ then
   Remove $b_{i}$ from $\mathcal{B}^{\prime}$. {Small-area pruning}
  end if
  if $\mathbf{IoU}\!\big(b_{i},\{\,b_{j}\in\mathcal{B}\mid\mathbf{colordiff}(b_{i},b_{j})<0.03\,\}\big)>0.9$ then
   Remove $b_{i}$ from $\mathcal{B}^{\prime}$. {Redundant-overlap pruning}
  end if
end for
Densification phase
$\mathcal{R}\leftarrow\mathbf{ConnectedComponents}(E)$.
$\mathcal{R}_{\mathbf{sorted}}\leftarrow\mathbf{SortByArea}(\mathcal{R})$.
for each region $r_{j}\in\mathcal{R}_{\mathbf{sorted}}$ do
  if curve budget allows then
   Insert a new closed Bézier curve into $r_{j}$.
  end if
end for
return $\mathcal{B}^{\prime}$.
Appendix C Convert Bézier Splatting to Standard SVG Format
As described in Algorithm 2, we convert our optimized Bézier Splatting representation into a standard SVG file for compatibility with downstream vector editing tools. Each curve is represented by a set of $3k+1$ control points $\mathcal{C}_{i}\in\mathbb{R}^{(3k+1)\times 2}$, corresponding to $k$ continuous cubic Bézier segments. These control points are normalized to the range [1,1][-1,1] and are first transformed into pixel coordinates according to the canvas size $(W,H)$.
@@ Appendix C Convert Bézier Splatting to Standard SVG Format
Algorithm 2 Convert Bézier Splatting to Standard SVG
@@
Appendix D Bézier Splatting Supports Flexible Curve Attributes
All 2D Gaussians in our method are generated via a differentiable sampling algorithm, which makes it straightforward to incorporate user-defined shape attributes, such as linear-gradient fills in color or opacity. To demonstrate the extensibility of our model, we evaluate the multi-opacity scheme for open curves: each Bézier segment within an open curve is assigned an independent opacity value. This design enables users to either maintain consistent sampling point appearance across segments or apply customized interpolation strategies. This strategy significantly improves vectorization quality for open curves, as shown in Table  7. Since closed curves are the primary representation format in SVG, we adopt a uniform setting for opacity and color in those regions to ensure a full compatibility with existing vector graphic tools. Nonetheless, users can still follow the above multi-opacity scheme to implement linear-gradient fills for both opacity and color, enhancing the flexibility and expressiveness of vector graphics.
Table 7: An ablation study on the number of opacity parameters per curve, evaluated on 512 open curves with DIV2K dataset [Timofte_2017_CVPR_Workshops].
@@ Appendix D Bézier Splatting Supports Flexible Curve Attributes
Appendix E More Comparisons
We present more qualitative comparisons on the DIV2K [Timofte_2017_CVPR_Workshops] dataset. As shown in Fig. 9, Fig. 10, Fig. 11, Fig. 12, and Fig. 13, our method shows consistent improvements over the baselines, including better preservation of fine details, higher rendering fidelity, fewer visual artifacts, and more accurate geometric structures. Furthermore, as the number of curves increases, our method demonstrates a significantly improved representational ability, enabling more precise reconstruction of complex shapes, and fine-grained structures.
We further evaluate our method on another natural image dataset, Kodak [kodak1999]. Due to the slow vectorization speed of LIVE [xu2022live], we compare with DiffVG on open curves only. As shown in Table 8, our method consistently outperforms DiffVG across quantitative metrics including PSNR, SSIM, and LPIPS.
Table 8: A quantitative evaluation on the Kodak dataset [kodak1999] with 256 to 1024 curves. We report results on open curves.
@@ Appendix E More Comparisons
Appendix F More Image Vectorization Results
Fig. 14, Fig. 15, and Fig. 16 show more results on natural images from DIV2K [Timofte_2017_CVPR_Workshops] and Kodak [kodak1999], as well as animation images from DanbooRegion [DanbooRegion2020], respectively. The results demonstrate that both the global structure and local texture of the images are well reconstructed, highlighting the effectiveness of our approach in capturing fine details and complex shapes.
Figure 9: A qualitative comparison of our method and the existing differentiable VG rasterization method on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 10: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 11: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 12: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 13: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 14: More image vectorization results by Bézier Splatting on DIV2K dataset [Timofte_2017_CVPR_Workshops].
Figure 15: More image vectorization results by Bézier Splatting on Kodak [kodak1999] dataset.
Figure 16: More image vectorization results by Bézier Splatting on DanbooRegion dataset [DanbooRegion2020].